OpenStudy (anonymous):

Determine the probability that at least one head appears in three flips of a fair coin. Round your answer to one decimal place.

OpenStudy (anonymous):

is it .875

OpenStudy (anonymous):

all 8 options are HHH,HHT,HTH, HTT, THH THT TTH TTT. (8 in total) seven do have atleast one head so 7/8.

OpenStudy (anonymous):

or u cud say that P(3 heads) = 1/8 P(at least one head) = 1- p(3 heads) = 1- 1/8 = 7/8

OpenStudy (anonymous):

= 0.875 = 0.9 (rounded off to one decimal place)